# Upper Approximation Bounds for Neural Oscillators

### v1

@@ 3.1 Causal Continuous Operators
$B_{\Omega}=2\max\left(TB_{K},B_{K},T\right)$, $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, $L_{{\psi_{j}}}\left(\cdot\right)$} is in Eq. (17), ${\phi_{\Phi}}(\cdot)$ is the monotonic modulus of continuity of $\Phi$ in Assumption 6, and $c_{\rho,2}$ is the independent constant in Lemma 1. $\it{\Gamma}(\cdot)$ has one hidden layer and the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}$, $w_{\it{\Gamma}}$, and $w_{{\it{\Gamma},{\mathrm{out}}}}$, respectively
@@
Remark 7: An upper approximation bound of ReLU MLPs by Yarotsky (2017) has been employed in Theorem 3 of Liu et al. (2024) to establish an upper generalization bound of neural operators. The result by Yarotsky (2017) is not employed in this study because its approximation bound constant depends on the argument dimension of a target function and this dependence is implicit. This implicit relationship hinders the establishment of the explicit relationship between the final approximation error $\varepsilon_{\bf{y}}$} in Eq. (19) and $M_{\it{\Gamma}}$ controlling the width of the hidden layer of the MLP ${\it{\Gamma}}(\cdot)$.
Remark 8: Theorem 1 indicates that the approximation error of the neural oscillator in Eq. (1) consists of two parts. The first one is from approximating $\mathbf{u}(t)$ using the sine transform and the second one is from approximating the continuous mapping $\Psi\left(\mathbf{x},\mathbf{u}_{0},t\right)$ in Lemma 3 using the MLP ${\it{\Pi}}(\cdot)$. Eq. (20) indicates that the decay rate $1/\left[p\left(M_{\it{\Gamma}}+1\right)+1\right]$ in the second error polynomially diminishes as the width $w_{\it{\Gamma}}=2pM_{\it{\Gamma}}$ of the hidden layer of the first MLP ${\it{\Gamma}}(\cdot)$ increases. Thus, the neural oscillator in Eq. (1) for approximating the causal continuous operators between continuous temporal function spaces also faces the curse of parametric complexity, which is a common problem in neural operators (Lanthaler, 2023; Lanthaler and Stuart, 2025). A recent result on the quantitative estimates of the number of parameters required in neural operators, presented in Proposition 1 of Kratsios et al. (2024), is similar to the second term on the right side of Eq. (20).
@@
One method to mitigate the curse of parametric complexity is applying other super-expressive activation functions (Maiorov and Pinkus, 1999; Shen et al., 2021; Yarotsky, 2021; Zhang et al., 2022). For instance, Theorem 4 in Maiorov and Pinkus (1999) demonstrates the existence of an activation function that enables a two-hidden-layer neural network architecture to approximate continuous functions with arbitrary accuracy, even when the widths of the hidden layers are held constant. Theorem 1.1 and Corollary 1.2 in Shen et al. (2021) provide a three-hidden-layer MLP architecture, called Floor-Exponential-Step (FLES) network, utilizing the floor function $\lfloor x\rfloor$, the exponential function $2^{x}$, and the step function as the activation function in each neuron. FLES is capable of approximating arbitrary continuous functions with the number of parameters scaling logarithmically with the reciprocal of the approximation error. Theorem 1 in Zhang et al. (2022) introduces a neural network architecture that employs an activation function combining a softsign activation function with a periodic triangular-wave function. This neural network architecture with a width $36d(2d+1)$ and a depth 11 can approximate any continuous function on a $d$-dimensional hypercube within an arbitrarily small error. Super-expressive activation functions have been utilized to mitigate the curse of parametric complexity in neural operators (Schwab et al., 2023). Since the explicit expressions for the Lipschitz constant $L_{\it{\psi}_{j}}(M)$ in Eq. (17) and the linear modulus of continuity $\phi_{\it{\psi}_{j},M}(x)$ in Eq. (18) of the continuous mapping $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$, along with its valid compact argument domain $\Omega_{M}\left(TB_{K},B_{K},T\right)$, have been derived, the super-expressive activation functions mentioned above can be directly applied to the MLP ${\it{\Pi}}(\cdot)$ to mitigate the curse of parametric complexity in the neural oscillator. On the other hand, Lanthaler (2024) has demonstrated that decreasing the required number of parameters in neural operators using super-expressive neural networks may come at the expense of requiring larger model parameter values.
@@ 3.2 Second-Order Dynamical Systems
In this subsection, an upper approximation bound of the neural oscillator in Eq. (1) for approximating uniformly asymptotically incrementally stable second-order dynamical systems is derived. First, the asymptotically incrementally stable second-order dynamical system is briefly introduced. Subsequently, Lemma 5 and Lemma 6 establish the relationships between the solution error of two ODEs subject to the same input functions and the error of the derivative functions of the two ODEs. Finally, by utilizing the result in Lemma 6 and the error bound for the MLP approximation of Barron class functions (Barron, 1993), the upper approximation bound of the neural oscillator is demonstrated in Theorem 2. This subsection builds on the methodology of Hanson and Raginsky (2020) and offers technical contributions through detailed mathematical derivations, as elaborated in the accompanying remarks. The proofs for all lemmas and Theorem 2 in this subsection are presented in B.
@@
$\begin{array}[]{l}{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})=\int_{0}^{t}{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}}\left\{{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right\}{\rm{d}}\tau}\end{array}$ (26)
@@
Remark 9: Being analogous to the fundamental Lemma 3.1.2 in Van Handel (2007), which has been employed by Hanson and Raginsky (2020) to derive the approximation bound of RNNs for modeling uniformly asymptotically incrementally stable first-order dynamical systems, Lemma 5 in this study provides a more rigorous proof. In the proof of the step involving Eqs. (81) and (82) in B.1, the proof of the original lemma in Van Handel (2007) utilizes ${\mathbf{z}}_{1}(\tau)={\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ but not ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)={\mathbf{z}}_{\tau}$ with $\hat{\tau}=\tau$ and replaces an arbitrary ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ with its corresponding ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ under the same $\mathbf{u}(t)$ to prove Eq. (82). However, this proof method is not rigorous. For an arbitrary time instant $\tau$ ($0\leq\tau\leq T$), the range of all possible values of ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ depends on the differential equation governed by $g_{1}(\cdot)$ and all $\mathbf{u}(t)\in K$. It is evident that an arbitrary value within the range of all possible values of ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ satisfies Eq. (81). However, it remains uncertain whether all possible values of ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ at the same time instant $\tau$ driven by $\mathbf{u}(t)\in K$ lie within the range of the values of ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$. If not, it is unclear whether ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ satisfies Eq. (81) for all $0\leq\tau\leq T$ and all $\mathbf{u}(t)\in K$. In this study, employing an arbitrary $\tau$ and an arbitrary ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)={\mathbf{z}}_{\tau}$ with $\hat{\tau}=\tau$ in Eq. (81) resolves this uncertainty. ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)$ is the solution at $^{\hat{\tau}}$ caused by the initial condition ${\mathbf{z}}_{\tau}$ at $\tau$. Being an initial condition, the value of ${\mathbf{z}}_{\tau}$ can be an arbitrary vector within $\Omega_{\mathbf{z}}$. ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)={\mathbf{z}}_{\tau}$ with $\hat{\tau}=\tau$ indicates that the value of ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)$ can also be an arbitrary vector within $\Omega_{\mathbf{z}}$. This implies that all vectors within $\Omega_{\mathbf{z}}$ satisfy Eq. (81). From the assumption of Lemma 5, the possible values of ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ fall within $\Omega_{\mathbf{z}}$ for all $\tau\in[0,T]$ and all $\mathbf{u}(t)\in K$. Consequently, all possible values of ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ satisfy Eq. (81).
@@
$\begin{array}[]{l}\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\leq{e^{T{L_{{g_{1}}}}}}T\mathop{\max}\limits_{\tau\in[0,T]}\left|{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right|\end{array}$ (27)
@@
$\begin{array}[]{l}\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right|_{2}}{B_{\beta}}\end{array}$ (28)
@@
Remark 10: As shown in D, the result in Eq. (27) can also be proven using the method based on the Grönwall’s inequality (Ames and Pachpatte, 1997), which was frequently employed in deriving approximation bounds of RNNs for modeling first-order dynamical systems (Funahashi and Nakamura, 1993; Chow and Li, 2000). The proof of Lemma 6 shows that the method based on Lemma 5, which enhances the proof of Lemma 3.1.2 in Van Handel (2007), can simultaneously obtain both categories of the bounds in Eq. (27) and Eq. (28) for the difference between the solutions of two different second-order ODEs subject to the same input functions. This advancement enhances the theoretical understanding of the second-order ODEs. Lemma 6 will be utilized in the subsequent Theorem 2 to establish the approximation bound of the neural oscillator for modeling uniformly asymptotically incrementally stable second-order dynamical systems.
@@
with the initial conditions ${\hat{\mathbf{x}}}(0)={{\hat{\mathbf{x}}}}^{\prime}(0)=\mathbf{0}$, where ${\hat{\mathbf{x}}}(t)=[{\hat{x}}_{1}(t),{\hat{x}}_{2}(t),\dots,$ ${\hat{x}}_{r}(t)]^{\top}$ and ${\hat{\mathbf{y}}}(t)=\left[{\hat{y}}_{1}(t),{\hat{y}}_{2}(t),\dots,{\hat{y}}_{q}(t)\right]^{\top}$. The functions $g(\cdot)=[g_{1}(\cdot),g_{2}(\cdot),$ $\dots,g_{r}(\cdot)]^{\top}$ and $h(\cdot)=\left[h_{1}(\cdot),h_{2}(\cdot),...h_{q}(\cdot)\right]^{\top}$ are Lipschitz continuous over ${\mathbb{R}}^{2r+p}$ and $\mathbb{R}^{r}$, respectively, with arbitrary finite-order derivatives. Assume that the dynamical system in Eq. (29) is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on a domain $\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}\subset\mathbb{R}^{2r+p}$, where $B_{{\mathbf{x}}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}}$, $B_{{\mathbf{x}}^{\prime}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}^{\prime}}$, $\alpha$ is a positive coefficient, $B_{\hat{\mathbf{x}}}$ and $B_{\hat{\mathbf{x}}^{\prime}}$ are the bounds of ${\hat{\mathbf{x}}}(t)$ and ${\hat{\mathbf{x}}}^{\prime}(t)$, respectively, $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, and $B_{\beta{g}}$ is the stability bound of the dynamical system in Eq. (25), then, for arbitrary positive errors $\varepsilon_{1}\leq\alpha/r$ and $\varepsilon_{2}$, there exist two one-hidden-layer MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ employing $\sigma_{\mathrm{ReLU}}(\cdot)$, such that for all $\mathbf{u}(t)\in K$, the corresponding solutions $\mathbf{y}(t)$ to Eq. (1) and ${\hat{\mathbf{y}}}(t)$ to Eq. (29) satisfy
@@
where $C_{\it{\Pi}}$ is a constant only depending on $h(\cdot)$, $\sigma_{\mathrm{ReLU}}(\cdot)$, $r$, and $B_{\mathbf{x}}$

### v2

@@ 3.1 Causal Continuous Operators
$B_{\Omega}=2\max\left(TB_{K},B_{K},T\right)$, $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, $L_{{\psi_{j}}}\left(\cdot\right)$} is in Eq. (17), ${\phi_{\Phi}}(\cdot)$ is the monotonic modulus of continuity of $\Phi$ in Assumption 6, and $c_{K}$ is the independent constant in Lemma 1. $\it{\Gamma}(\cdot)$ has one hidden layer and the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}$, $w_{\it{\Gamma}}$, and $w_{{\it{\Gamma},{\mathrm{out}}}}$, respectively
@@
Remark 7: The approximation error of the neural oscillator in Theorem 1, which can be expressed as ${\varepsilon_{\bf{y}}}={\phi_{\Phi}}\left\{O\left[\left(\ln{M_{{\it{\Gamma}}}}\right)^{2}/{M_{\it{\Gamma}}}\right]\right\}+O\left\{M_{\it{\Gamma}}^{5}H_{\it{\Pi}}^{-1/\left[p\left(M_{{\it{\Gamma}}}+1\right)+1\right]}\right\}$ in Eq (20), consists of two parts. The second one arises from approximating the continuous mapping $\Psi\left(\mathbf{x},\mathbf{u}_{0},t\right)$ in Lemma 3 using the ReLU MLP ${\it{\Pi}}(\cdot)$ and it is derived from Theorem 1 by Hanin (2019), where $p\left(M_{{\it{\Gamma}}}+1\right)+1$ in $\varepsilon_{\bf{y}}$} is the argument dimension of $\Psi({\bf{x}},{{\bf{u}}_{0}},t)$. Theorem 4 in Yarotsky (2017) provides a lower bound with a convergence rate of 2/d-2/d for approximating Lipschitz continuous functions using ReLU MLPs, where $d$ is the argument dimension. A corresponding upper bound $O\left[\left(\ln{W}/W\right)^{2/d}\right]$, where $W$ represents the total parameter number of the employed ReLU MLP, was established in Theorem 4.1 of Yarotsky and Zhevnerchuk (2020). Although the upper approximation bound by Yarotsky and Zhevnerchuk (2020) is closer to the lower bound with a convergence rate of 2/d-2/d than 1/d-1/d of Theorem 1 by Hanin (2019), its approximation bound constant depends on the argument dimension of a target function, and this dependence is implicit. This implicit relationship hinders the establishment of the explicit relationship between the final approximation error $\varepsilon_{\bf{y}}$} in Eq. (19) and $M_{\it{\Gamma}}$ controlling the width of the hidden layer of the MLP ${\it{\Gamma}}(\cdot)$. The convergence rate of an upper bound for neural operators presented in Proposition 1 of Kratsios et al. (2024) is also in the form of 1/d-1/d. The first part of $\varepsilon_{\bf{y}}$} comes from approximating the input function $\mathbf{u}(t)$ using the sine transform. The convergence rate of a lower bound for approximating compactly supported continuous functions via the Fourier series is 1-1 (DeVore, 1998). Compared to this lower bound, the additional logarithmic term $\left(\ln{M_{{\it{\Gamma}}}}\right)^{2}$ in Theorem 1 is due to the discontinuities of $\tilde{\mathbf{f}}_{t}(\tau)$ in Eq. (44) of A.1, as discussed in Remark 2. The sine transform encoding method is linear, whereas the neural ODE in the neural oscillator can represent a wide range of nonlinear dynamical systems. Whether there exists an encoding method based on nonlinear neural ODEs that can achieve a faster bound than $O\left[\left(\ln{M_{{\it{\Gamma}}}}\right)^{2}/{M_{\it{\Gamma}}}\right]$ in Theorem 1 remains an interesting direction for future investigation.
Remark 8: Theorem 1 in Zhang et al. (2024) demonstrates that for a large set of activation functions, including the commonly used tanh function $\sigma_{\text{tanh}}(x)=\left(e^{x}-e^{-x}\right)/\left(e^{x}+e^{-x}\right)$ and sigmoid function $\sigma_{\text{sigmoid}}(x)=1/\left(1+e^{-x}\right)$, any ReLU MLP with the maximum width $W$ and depth $H$ can be approximated to arbitrary accuracy by an MLP using any activation function from the activation function set, with the maximum width $3W$ and depth $2H$. Furthermore, Lemma 3.7 in Lanthaler et al. (2023) shows that the sine transform $\mathcal{L}_{t}\mathbf{u}(\omega)$ in Eq. (10) can be approximated to arbitrary accuracy by the solution of a second-order neural ODE with an activation function $\sigma(x)$ satisfying $\sigma(0)=0$ and $\sigma^{\prime}(0)=1$. These two conditions are satisfied by $\sigma_{\text{tanh}}(x)$ and the scaled sigmoid function $2\sigma_{\text{sigmoid}}(2x)-1$. Therefore, approximation bounds with slightly larger bound constants, that are similar to the result in Theorem 1, can be achieved when $\sigma_{\text{tanh}}(x)$ or $\sigma_{\text{sigmoid}}(x)$ is utilized in the neural oscillator.
Remark 9: Theorem 1 first decomposes the input $\mathbf{u}(t)$ into $M_{{\it{\Gamma}}}$ components, and subsequently maps these $M_{{\it{\Gamma}}}$ components to the output $\mathbf{y}(t)$. Since the function set $K$ containing $\mathbf{u}(t)$ can be infinite-dimensional, $M_{{\it{\Gamma}}}$ must tend to infinity for the first part of $\varepsilon_{\bf{y}}$} in Eq. (20) to vanish. However, this will cause the convergence rate $1/\left[p\left(M_{\it{\Gamma}}+1\right)+1\right]$ in the second part of $\varepsilon_{\bf{y}}$} to diminish towards zero. Thus, the neural oscillator in Eq. (1) for approximating the causal continuous operators between continuous temporal function spaces also faces the curse of parametric complexity, which is a common problem in neural operators (Lanthaler, 2023; Lanthaler and Stuart, 2025).
@@
One method to overcome the curse of parametric complexity is applying other super-expressive activation functions (Maiorov and Pinkus, 1999; Shen et al., 2021; Yarotsky, 2021; Zhang et al., 2022). For instance, Theorem 4 in Maiorov and Pinkus (1999) demonstrates the existence of an activation function that enables a two-hidden-layer neural network architecture to approximate continuous functions with arbitrary accuracy, even when the widths of the hidden layers are held constant. Theorem 1.1 and Corollary 1.2 in Shen et al. (2021) provide a three-hidden-layer MLP architecture, called Floor-Exponential-Step (FLES) network, utilizing the floor function $\lfloor x\rfloor$, the exponential function $2^{x}$, and the step function as the activation function in each neuron. FLES is capable of approximating arbitrary continuous functions with the number of parameters scaling logarithmically with the reciprocal of the approximation error. Theorem 1 in Zhang et al. (2022) introduces a neural network architecture that employs an activation function combining a softsign activation function with a periodic triangular-wave function. This neural network architecture with a width $36d(2d+1)$ and a depth 11 can approximate any continuous function on a $d$-dimensional hypercube within an arbitrarily small error. Super-expressive activation functions have been utilized to overcome the curse of parametric complexity in neural operators (Schwab et al., 2023). Since the explicit expressions for the Lipschitz constant $L_{\it{\psi}_{j}}(M)$ in Eq. (17) and the linear modulus of continuity $\phi_{\it{\psi}_{j},M}(x)$ in Eq. (18) of the continuous mapping $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$, along with its valid compact argument domain $\Omega_{M}\left(TB_{K},B_{K},T\right)$, have been derived, the super-expressive activation functions mentioned above can be directly applied to the MLP ${\it{\Pi}}(\cdot)$ to overcome the curse of parametric complexity in the neural oscillator. On the other hand, Lanthaler (2024) has demonstrated that decreasing the required number of parameters in neural operators using super-expressive neural networks may come at the expense of requiring larger model parameter values.
@@ 3.2 Second-Order Dynamical Systems
In this subsection, an upper approximation bound of the neural oscillator in Eq. (1) for approximating uniformly asymptotically incrementally stable second-order dynamical systems is derived. First, the asymptotically incrementally stable second-order dynamical system is briefly introduced. Subsequently, Lemma 5 and Lemma 6 establish the relationships between the solution error of two ODEs subject to the same input functions and the error of the derivative functions of the two ODEs. Finally, by utilizing the result in Lemma 6 and the error bound for the MLP approximation of Barron class functions (Barron, 1993; Siegel and Xu, 2023), the upper approximation bound of the neural oscillator is demonstrated in Theorem 2. This subsection builds on the methodology of Hanson and Raginsky (2020) and offers technical contributions through detailed mathematical derivations, as elaborated in the accompanying remarks. The proofs for all lemmas and Theorem 2 in this subsection are presented in B.
@@
${{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})=\int_{0}^{t}{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}}\left\{{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right\}{\rm{d}}\tau}$ (26)
@@
Remark 10: Being analogous to the fundamental Lemma 3.1.2 in Van Handel (2007), which has been employed by Hanson and Raginsky (2020) to derive the approximation bound of RNNs for modeling uniformly asymptotically incrementally stable first-order dynamical systems, Lemma 5 in this study provides a more rigorous proof. In the proof of the step involving Eqs. (86) and (87) in B.1, the proof of the original lemma in Van Handel (2007) utilizes ${\mathbf{z}}_{1}(\tau)={\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ but not ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)={\mathbf{z}}_{\tau}$ with $\hat{\tau}=\tau$ and replaces an arbitrary ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ with its corresponding ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ under the same $\mathbf{u}(t)$ to prove Eq. (87). However, this proof method is not rigorous. For an arbitrary time instant $\tau$ ($0\leq\tau\leq T$), the range of all possible values of ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ depends on the differential equation governed by $g_{1}(\cdot)$ and all $\mathbf{u}(t)\in K$. It is evident that an arbitrary value within the range of all possible values of ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ satisfies Eq. (86). However, it remains uncertain whether all possible values of ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ at the same time instant $\tau$ driven by $\mathbf{u}(t)\in K$ lie within the same range of the values of ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$. If not, it is unclear whether ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ satisfies Eq. (86) for all $0\leq\tau\leq T$ and all $\mathbf{u}(t)\in K$. In this study, employing an arbitrary $\tau$ and an arbitrary ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)={\mathbf{z}}_{\tau}$ with $\hat{\tau}=\tau$ in Eq. (86) resolves this uncertainty. ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)$ is the solution at $^{\hat{\tau}}$ caused by the initial condition ${\mathbf{z}}_{\tau}$ at $\tau$. Being an initial condition, the value of ${\mathbf{z}}_{\tau}$ can be an arbitrary vector within $\Omega_{\mathbf{z}}$. ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)={\mathbf{z}}_{\tau}$ with $\hat{\tau}=\tau$ indicates that the value of ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)$ can also be an arbitrary vector within $\Omega_{\mathbf{z}}$. This implies that all vectors within $\Omega_{\mathbf{z}}$ satisfy Eq. (86). From the assumption of Lemma 5, the possible values of ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ fall within $\Omega_{\mathbf{z}}$ for all $\tau\in[0,T]$ and all $\mathbf{u}(t)\in K$. Consequently, all possible values of ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ satisfy Eq. (86).
@@
$\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\leq{e^{T{L_{{g_{1}}}}}}T\mathop{\max}\limits_{\tau\in[0,T]}\left|{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right|$ (27)
@@
$\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right|_{2}}{B_{\beta}}$ (28)
@@
Remark 11: As shown in D, the result in Eq. (27) can also be proven using the method based on the Grönwall’s inequality (Ames and Pachpatte, 1997), which was frequently employed in deriving approximation bounds of RNNs for modeling first-order dynamical systems (Funahashi and Nakamura, 1993; Chow and Li, 2000). The proof of Lemma 6 shows that the method based on Lemma 5, which enhances the proof of Lemma 3.1.2 in Van Handel (2007), can simultaneously obtain both categories of the bounds in Eqs. (27) and (28) for the difference between the solutions of two different second-order ODEs subject to the same input functions. This advancement enhances the theoretical understanding of the second-order ODEs. Lemma 6 will be utilized in the subsequent Theorem 2 to establish the approximation bound of the neural oscillator for modeling uniformly asymptotically incrementally stable second-order dynamical systems.
@@
with the initial conditions ${\hat{\mathbf{x}}}(0)={{\hat{\mathbf{x}}}}^{\prime}(0)=\mathbf{0}$, where ${\hat{\mathbf{x}}}(t)=\left[{\hat{x}}_{1}(t),{\hat{x}}_{2}(t),\dots,{\hat{x}}_{r}(t)\right]^{\top}$ and ${\hat{\mathbf{y}}}(t)=\left[{\hat{y}}_{1}(t),{\hat{y}}_{2}(t),\dots,{\hat{y}}_{q}(t)\right]^{\top}$. $g(\cdot)=\left[g_{1}(\cdot),g_{2}(\cdot),\dots,g_{r}(\cdot)\right]^{\top}$ satisfies that for every compact domain $\Omega\subset$ ${\mathbb{R}}^{2r+p}$, there exists a Barron function extension $\tilde{g}_{\Omega}(\cdot)=g(\cdot)$ on $\Omega$ and the spectral Barron norms of all extensions are uniformly bounded. $h(\cdot)=\left[h_{1}(\cdot),h_{2}(\cdot),\dots,h_{q}(\cdot)\right]^{\top}$ satisfies that for every compact domain $\Omega\subset$ ${\mathbb{R}}^{r}$, there exists a Barron function extension $\tilde{h}_{\Omega}(\cdot)=h(\cdot)$ on $\Omega$ and the spectral Barron norms of all extensions are uniformly bounded. Assume that the dynamical system in Eq. (29) is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on a domain $\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}\subset\mathbb{R}^{2r+p}$, where $B_{{\mathbf{x}}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}}$, $B_{{\mathbf{x}}^{\prime}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}^{\prime}}$, $\alpha$ is a positive coefficient, $B_{\hat{\mathbf{x}}}$ and $B_{\hat{\mathbf{x}}^{\prime}}$ are the bounds of ${\hat{\mathbf{x}}}(t)$ and ${\hat{\mathbf{x}}}^{\prime}(t)$, respectively, $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, and $B_{\beta{g}}$ is the stability bound of the dynamical system in Eq. (25), then, for arbitrary positive errors $\varepsilon_{1}\leq\alpha/r$ and $\varepsilon_{2}$, there exist two one-hidden-layer MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ employing $\sigma_{\mathrm{ReLU}}(\cdot)$, such that for all $\mathbf{u}(t)\in K$, the corresponding solutions $\mathbf{y}(t)$ to Eq. (1) and ${\hat{\mathbf{y}}}(t)$ to Eq. (29) satisfy
@@
where $C_{\it{\Pi}}$ is a constant only depending on $h(\cdot)$, $\sigma_{\mathrm{ReLU}}(\cdot)$, $r$, and $B_{\mathbf{x}}$.